# Robust Low-Cost Drone Detection and Classification Using Convolutional Neural Networks in Low SNR Environments

### v2

@@ Abstract
I Introduction
@@ II-B Data Preparation
The selected drone signal vectors with were normalised to a carrier power of per sample, i.e. only the part of the signal vector containing drone bursts was considered for the power calculation ( samples out of ). This was achieved by identifying the bursts as those samples where a smoothed energy was above a threshold. The signal vectors are thus normalised by
@@ II-C Development Dataset
As described in Sec. II-B, the drone signals were mixed with noise. More specifically, of the drone signals were mixed with Labnoise (Bluetooth + Wi-Fi + Amplifier) and with Gaussian noise. In addition, we created a separate noise class by mixing Labnoise and Gaussian noise in all possible combinations (i.e., Labnoise Labnoise, Labnoise Gaussian noise, Gaussian noise Labnoise, and Gaussian noise Gaussian noise). For the drone signal classes, as for the noise class, the number of samples for each SNR level was evenly distributed over the interval of SNRs dB in steps of dB, i.e., - samples per SNR level. The resulting number of samples per class is given in Tab. III.
@@ III-A Model Architecture and Training
As in [20] we chose the Visual Geometry Group (VGG) CNN architecture [24]. The main idea of this architecture is to use multiple layers of small () convolutional filters instead of larger ones. This is intended to increase the depth and expressiveness of the network, while reducing the number of parameters. There are several variants of this architecture, which differ in the number of convolutional layers ( and , respectively). We used a variant with a batch normalisation [25] layer after the convolutions, denoted as VGG11_BN to VGG19_BN. For the dense classification layer, we used linear units followed by linear units at the output (one unit per class).
@@ IV-A Classification performance in the Development Dataset
Figure 6 shows the resulting -fold mean balanced accuracy over SNRs dB in dB steps. Note that we do not show the standard deviation to keep the plot readable. In general, we observe a drastic degradation in performance from dB down to near chance level at dB.
@@ IV-B Embedding Space Visualisation
Figure 8 shows the 2D t-SNE visualisation of the VGG11_BN embeddings of test samples from the development dataset. It can be seen that each class forms a separate cluster. While the different drone signal clusters are rather small and dense, the noise cluster takes up most of the embedding space and even forms several sub-clusters. This is most likely due to the variety of the signals used in the noise class, i.e. Bluetooth and Wi-Fi signals plus Gaussian noise.
@@ V Discussion
Based on our experience and results, we see the following limitations of our work. The field test showed that the models can be used and work reliably (cf. Tab. LABEL:tab:field_test_acc_type_direction). However, it is the nature of a field test that the level of interference from WiFi/Bluetooth noise and the possible presence of other drones cannot be fully controlled. Furthermore, due to the limited space/distance between the transmitter and receiver in our field test setup, we were not able to clearly demonstrate the effect of free space attenuation on detection performance (cf. Tab. LABEL:tab:field_test_acc_distance).
@@
We have only tested a limited set of VGG architectures. It remains to be seen whether more recent architectures, such as the pre-trained Vision Transformer [32], generalise as well or better. We hope that our development dataset will inspire others to further optimise the model side of the problem and perhaps find a model architecture with better performance.
Another issue to consider is the occurrence of unknown drones, i.e. drones that are not part of the train set. Examining the embedding space (cf. IV-B) gives a first idea of whether a signal is clearly part of a known dense drone cluster or rather falls into the larger, less dense, noise cluster. We believe that a combination of an unsupervised deep autoencoder approach [33, 34] with an additional classification part (cf. [35]) would allow, first, to provide a stable classification of known samples and, second, to indicate whether a sample is known or rather an anomaly.

### v3

@@ Abstract
Index Terms:
Drone detection, UAV classification, Low signal-to-noise ratio, Robustness, Real-world field test.I Introduction
@@ II-B Data Preparation
The selected drone signal vectors with were normalised to a carrier power of per sample, i.e. only the part of the signal vector containing drone bursts was considered for the power calculation ( samples out of ). This was achieved by identifying the bursts as those samples where a smoothed energy was above a threshold.
The signal vectors are thus normalised by
@@ II-C Development Dataset
As described in Sec. II-B, the drone signals were mixed with noise (cf. Eqs. -). More specifically, of the drone signals were mixed with Labnoise (Bluetooth + Wi-Fi + Amplifier) and with Gaussian noise. In the same way we created a single noise class by mixing Labnoise and Gaussian noise in all possible combinations (i.e., Labnoise Labnoise, Labnoise Gaussian noise, Gaussian noise Labnoise, and Gaussian noise Gaussian noise). This mixing was done as for the drone signals using Eq. . For instance, in case of a Labnoise Gaussian noise mix, refers to a Labnoise vector and to a generated Gaussian noise vector.
For the drone signal classes, as for the noise class, the number of samples for each SNR level was evenly distributed over the interval of SNRs dB in steps of dB, i.e., - samples per SNR level. The resulting number of samples per class is given in Tab. III.
@@ III-A Model Architecture and Training
As in [20] we chose the Visual Geometry Group (VGG) CNN architecture [24] due to its performance history across a wide range of tasks, including medical image classification [25], as backbone for object detection [26], and audio signal classification [27].
The main idea of this architecture is to use multiple layers of small () convolutional filters instead of larger ones. This is intended to increase the depth and expressiveness of the network, while reducing the number of parameters. There are several variants of this architecture, which differ in the number of convolutional layers ( and , respectively). We used a variant with a batch normalisation [28] layer after the convolutions, denoted as VGG11_BN to VGG19_BN. For the dense classification layer, we used linear units followed by linear units at the output (one unit per class).
@@
Model training was performed for epochs with a batch size of . The PyTorch [29] implementation of the Adam algorithm [30] was used with a learning rate of , betas and weight decay of . It was conducted on a NVIDIA A100 GB GPU, requiring GB of memory and h of computational time for the VGG11_BN training, and GB of memory and h of computational time for the VGG19_BN.
@@ IV-A Classification performance in the Development Dataset
Figure 6 shows the resulting -fold mean balanced accuracy over SNRs dB in dB steps. Note that we do not show the results for SNRs dB, as those are simply saturated at 100 balanced accuracy.
In general, we observe a drastic degradation in performance from dB down to near chance level at dB.
@@ IV-B Embedding Space Visualisation
Figure 8 shows the 2D t-SNE visualisation of the VGG11_BN embeddings of test samples from the development dataset. Each class forms a separate cluster. While the different drone signal clusters are rather small and dense, the noise cluster takes up most of the embedding space and even forms several sub-clusters. This is most likely due to the variety of the signals used in the noise class, i.e. Bluetooth and Wi-Fi signals plus Gaussian noise. It can also be seen that the DJI and FutabaT14 classes are more difficult to separate from noise than other classes (cf. Fig. 7) as they tend to overlap at the clusters’ edges.
@@ V Discussion
Based on our experience and results, we see the following limitations of our work. The RF signals of drones were recorded in a controlled laboratory environment and augmented with WiFi/Bluetooth/Gaussian noise (cf. Sec. II). While the laboratory environment was necessary for the initial development and evaluation of the model, these signals do not fully replicate the complexities encountered in real-world scenarios. None the less, the field test showed that the models can be used and work reliably (cf. Tab. LABEL:tab:field_test_acc_type_direction) in real world conditions. However, it is the nature of a field test that the level of interference from WiFi/Bluetooth noise and the possible presence of other drones cannot be fully controlled. Furthermore, due to the limited space/distance between the transmitter and receiver in our field test setup, we were not able to clearly demonstrate the effect of free space attenuation on detection performance (cf. Tab. LABEL:tab:field_test_acc_distance).
@@
We have only tested a limited set of VGG architectures. It remains to be seen whether more recent architectures, such as the pre-trained Vision Transformer [35], generalise as well or better. Further, with the deployment of the model on limited computational resources in mind, the use of more efficient and less complex models, such as GoogLeNet [36] and ShuffleNet [37] should be investigated. Research could aim for a balance between detection accuracy and the computational demands of real-time processing. By doing so, the deployment of a robust, low-latency drone detection system in dynamic and resource-constrained settings could become a practical reality.
Another issue to consider is the occurrence of unknown drones, i.e. drones that are not part of the train set. Examining the embedding space (cf. IV-B) gives a first idea of whether a signal is clearly part of a known dense drone cluster or rather falls into the larger, less dense, noise cluster. We believe that a combination of an unsupervised deep autoencoder approach [38, 39] with an additional classification part (cf. [40]) would allow, first, to provide a stable classification of known samples and, second, to indicate whether a sample is known or rather an anomaly. This could be accompanied by additional field test to further access the model’s generalisation ability to drones it was not trained on.
Finally, the current lack of standardised benchmarks makes it challenging to compare the performance of various models and systems objectively. We hope that our development dataset will inspire others to further optimise the model side of the problem and perhaps find a model architecture with better performance and/or efficiency.